A family determined the average cost of maintaining and operating the family car to be about $.30 per mile. On one trip, the family drove at an average rate of 50 miles per hour for a total of 6.5 hours. On the second trip, they drove at an average rate of 55 miles per hour for a total of 6 hours. which trip cost more? how much more?
Answer by richwmiller(17219)

50*6.5=325
55*6=330
330-325=5 miles more
5*.30=1.50
The second trip cost $1.50 more.